In 2005, for each of the 14 million people present in a country, 0.028 were born and 0.008 died during the year. Using exponential equation, the number of people present in 2015 is predicted as:
a. 25 millions
b. 17 millions
c. 20 millions
d. 18 millions

Grade:12

4 Answers

Akshat Jain
8 years ago
dN/dT = rN
 
\DeltaN/\DeltaT = rN
 
as we know that r = b-d
r =  0.028-0.008
  = 0.020
N = initial poulation = 14 million
 
\DeltaN/\DeltaT = 0.020x14 million
 
\DeltaN/10 = 0.28
\DeltaN = 2.8 million
\DeltaN = N- Ni = 2.8
Nf = 2.8 + 14 = 16.8 million
So answer is (b) 17 million.
 
Rinku
8 years ago
As given in question we are to use exponential equation.
Why we are using this equation to solve this problem ?
 Please tell me in detail ..
Akshat Jain
8 years ago
Exponential equation is integral form of this equation i.e. Nt =N0 ert
If you use this equation then you will fell difficulties to solve exponetial function.
But this equation can also be used to solve this problem..
Thanks For Such A Good Question ..
G Maheshbabu
7 years ago
 
dN/dT = rN
 
\DeltaN/\DeltaT = rN
 
as we know that r = b-d
r =  0.028-0.008
  = 0.020
N = initial poulation = 14 million
 
\DeltaN/\DeltaT = 0.020x14 million
 
\DeltaN/10 = 0.28
\DeltaN = 2.8 million
\DeltaN = N- Ni = 2.8
Nf = 2.8 + 14 = 16.8 million
So answer is (b) 17 million.